# Transient dynamical phase diagram of the spin–boson model

### v1

@@ Abstract
We investigate the real-time dynamics of the sub-Ohmic spin–boson model across a broad range of coupling strengths, using the numerically exact inchworm quantum Monte Carlo algorithm. From short- and intermediate-time dynamics starting from an initially decoupled state, we extract signatures of the zero-temperature quantum phase transition between localized and delocalized states. We show that the dynamical phase diagram thus obtained differs from the equilibrium phase diagram in both the values of critical couplings and the associated critical exponents. We also identify and quantitatively analyze two competing mechanisms for the crossover between coherent oscillations and incoherent decay. Deep in the sub-Ohmic regime, the crossover is driven by the damping of the oscillation amplitude, while closer to the Ohmic regime the oscillation frequency itself drops sharply to zero at large coupling.
@@ Introduction.
Although at finite $\omega_{c}$ delocalization is possible even at zero temperature, it can take a very long time for the system to delocalize [1]. Studies of the QPT have thus so far been performed directly in equilibrium, even though experimental probing of the dynamics can only access short- and intermediate-time properties. Another important feature manifested by the transient dynamics is the coherence of the decay process [46, 47, 35, 48, 49]. Usually, the change from well-defined oscillations at weak spin–bath coupling to incoherent decay at strong coupling is characterized as a crossover.
In this Letter, we extract signatures of the localization transition directly from the dynamics, showing that the result differs dramatically from the equilibrium one for $s\gtrsim 0.4$. We also reveal two distinct mechanisms that drive the change from coherent to incoherent decay: Overdamping of the oscillation amplitude that occurs at all values $10<s<$ 1; and a sharp decrease in oscillation frequency that is only observed for $s\gtrsim 0.5$. While the former is indeed a smooth crossover, the latter mechanism has the hallmarks of a sudden transition.
@@
The transient behavior of the sub-Ohmic spin–boson model depends strongly on the bath initial condition [46]. One usually considers the “decoupled” initial condition (where the bath is decoupled from the spin subsystem) or the “shifted” initial condition (where the bath is at equilibrium with the spin subsystem state fixed) [49]. Here, the initial system is assumed to be decoupled and the total density matrix can be represented by the factorized state $\rho_{0}=\rho_{s}\otimes\rho_{b}$, where the spin subsystem is in state $1$ ($\rho_{s}=|1\rangle\langle 1|$) and the bath in thermal equilibrium, $\rho_{b}=e^{-\beta H_{b}}/\textnormal{Tr}\left\{e^{-\beta H_{b}}\right\}$. Following the initial preparation, we focus on the dynamics of the population difference between the two spin states, $\langle\hat{\sigma}_{z}(t)\rangle=\textnormal{Tr}\left\{\rho_{0}e^{iHt}\hat{\sigma}_{z}e^{-iHt}\right\}$.
@@
To obtain numerically exact dynamics of $\langle\hat{\sigma}_{z}(t)\rangle$, we employ a real-time variant [50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59] of the continuous-time QMC algorithm [60, 61, 62]. To bypass the dynamical sign problem, which makes it exponentially difficult to access long times, we implemented an “inchworm” algorithm [40]. The idea behind this algorithm is that short-time propagators are less expensive to calculate than long-time ones, and can be recycled to construct easier expansions for propagators over longer timescales in subsequent Monte Carlo steps. The inchworm algorithm has been successfully applied to study the dynamics of the spin–boson model with the Debye spectral density [41, 42], in which two types of diagrammatic expansions were developed: the spin–bath coupling expansion and the diabatic coupling expansion (the latter is combined with a cumulant inchworm expansion). The results are consistent, although in different parameter regimes one of the expansions may be more efficient than the other [42]. In the context of the Ohmic/sub-Ohmic spin–boson model, we validated the accuracy of our inchworm Monte Carlo results by detailed comparisons with numerically exact results that are available at zero temperature using ML-MCTDH [63], see Supplemental Material (SM) [64] for more details.
@@ Results.
We only consider unbiased systems, i.e., $\epsilon=0$. We set $\omega_{c}=10\Delta$ throughout and use $\Delta$ as the unit of frequency. We have found inverse temperatures in the range $\beta\Delta\sim 20-100$ to be indistinguishable from the zero temperature limit within the statistical errors of the simulation.
In Fig. 1 we qualitatively compare the time-evolution of $\langle\sigma_{z}(t)\rangle$ as a function of coupling $\alpha$ in the deep sub-Ohmic ($s=0.2$) and near-Ohmic ($s=0.8$) regimes. In both regimes, the system undergoes a transition from a delocalized state at weak coupling to a localized state at strong coupling. The corresponding critical coupling $\alpha^{*}$, as obtained by criterion (a), increases with increasing $s$. We also observe the crossover between coherent oscillations at weak coupling and incoherent decay at strong coupling. In the deep sub-Ohmic regime, the crossover is damping-driven as the oscillation frequency remains essentially unchanged and the amplitude is more strongly damped when $\alpha$ increases (criterion (c)). In the near-Ohmic regime on the other hand, in addition to the decrease in amplitude, the oscillation frequency also decreases with increasing $\alpha$ and completely vanishes at large $\alpha$ (criterion (b)). The $\alpha$ dependence of the renormalized oscillation frequency can be captured qualitatively by a theoretical treatment based on the analytical renormalization group method used in Ref. [65]; the result is displayed as the dashed black curves representing oscillation peaks in the figure (see SM [64] for further discussion).
A quantitative analysis of the fit parameters obtained by fitting the numerical data to Eq. (4) sheds light on the nature of the localization/delocalization transition and the coherence/incoherence crossover. To distinguish between the localized and delocalized phase we examine the offset coefficient $c$ of the fit, shown in Fig. 2. A sharp transition from zero to finite $c$ is observed for $s\lesssim 0.5$. For larger $s$ (shown in Fig. 9 of the SM [64]), the offset $c$ increases more smoothly with $\alpha$, resulting in a larger uncertainty on the critical value $\alpha^{*}$. For the equilibrium QPT the critical exponents of the transition are known [2, 13, 14, 15]. At $s\leq 0.5$ the exponent $\beta$ defined via $\langle\sigma_{z}\rangle\propto(\alpha-\alpha_{c})^{\beta}$ equals $1/2$. In contrast, the corresponding scaling of the transient phase diagram analyzed in this work indicates a scaling exponent above 1, see SM [64] for more information.
@@
Next, to characterize the coherence/incoherence crossover, we focus on the $\alpha$-dependence of the renormalized frequency $\Omega/\Delta$, criterion (b), and its ratio to the damping coefficient $\Omega/\gamma_{1}$, criterion (c), both shown in Fig. 3. Interestingly, after a small initial dip, the frequency increases with increasing $\alpha$ for small values of $s$ ($s<0.5$). In contrast, for $s\gtrsim 0.5$ the frequency monotonically decreases with $\alpha$ and eventually sharply drops to zero at sufficiently strong coupling. This sharp frequency drop is not predicted by the renormalization group treatment [65, 64] and is distinct from the localization/delocalization transition. The widely accepted notion is that the crossover between coherent and incoherent dynamics is smooth. As opposed to the sharp frequency-driven incoherence transition, criterion (c) identifies the standard hallmark of a smooth incoherence crossover, which is driven by the over-damping of the oscillation. The $\Omega/\gamma_{1}$ ratio decreases with increasing $\alpha$ for all values of $s$ and we choose the (arbitrary) threshold $\Omega/\gamma_{1}=1$ to indicate the over-damping crossover. The two incoherence mechanisms are clearly distinct from each other, both in terms of their sharpness and in terms of the parameter range in which they occur.
@@
Figure 3: Two distinct hallmarks of the coherence/incoherence crossover: the renormalized oscillation frequency, $\Omega/\Delta$, (top panel) and the ratio between $\Omega$ and the corresponding damping coefficient $\gamma_{1}$ (bottom panel). The horizontal line in the bottom panel marks the $\Omega/\gamma_{1}=1$ threshold.
@@
A summary of our quantitative results is presented through the phase diagram in the parameter space of $\alpha$ and $s$, see Fig. 4. The transient critical coupling $\alpha^{*}$ distinguishing the localized and delocalized phases agrees well with the $\alpha_{c}$ of the equilibrium QPT for a model with sudden frequency cutoff [66, 20, 67] for sufficiently small $s$, where the critical couplings are also well-described by the analytical prediction from the generalized polaron ansatz [68]. For $s\gtrsim 0.4$ the two values begin to deviate. The frequency- and the damping-driven crossover lines depicted in Fig. 4 are clearly distinct from each other, with the frequency-driven crossover only occurring in the near-Ohmic regime, while the damping-driven crossover can occur at any value of $s$ (though it is numerically challenging to observe at small $s$). Coherent and incoherent regimes are observed on either side of the localization/delocalization line. For the Ohmic spin–boson model the location of the incoherence transition is known analytically to occur at $\alpha=0.5$ (Toulouse point) [2]. Our numerical results for the frequency-driven transition (criterion (b)) are consistent with the Toulouse point. The $\Omega=\gamma_{1}$ line of the damping-driven criterion (c) yields a different value for the Ohmic transition point.
@@
Figure 4: Transient dynamical phase diagram of the sub-Ohmic spin–boson model. The red shaded region is our numerical result for $\alpha^{*}$ that characterizes the dynamical localization-delocalization transition (criterion (a)), with the area marking a confidence interval. For $s\lesssim 0.4$ it agrees well with the equilibrium $\alpha_{c}$ from a QMC calculation for a model with sudden cutoff [20] (red circles and solid line) [69]. The dotted and dashed lines are the two characterizations of the coherent to incoherent , corresponding to criteria (b) and (c), respectively. The system tends towards incoherence above the lines, i.e. at higher values of $\alpha$. The frequency-driven crossover (blue dotted line) was only observed in the near-Ohmic regime, for the range of parameters considered. The damping-driven crossover (green dashed line), which we defined as $\gamma_{1}=\Omega$, was observed in a wider parameter range.
@@ Conclusions.
The inchworm algorithm used here is efficient over a wide range of parameters. Looking forward, it can also be used to explore dynamics and full charge/energy counting statistics at higher temperatures [70, 71, 72, 73], in more general models [74], or in the presence of intrinsically nonequilibrium drives such as multiple baths at different thermodynamic parameters. Extensions exist for investigating dynamics over very long timescales [75] and nonequilibrium steady states [76, 77]. On a more general note, our work points the way towards the ability to investigate and potentially control transient “phase diagrams” in numerous scenarios, opening up intriguing prospects for revealing new physics in experiments, theory and simulations.
Acknowledgements.
We thank Jianshu Cao, Jan von Delft, Carlos González-Gutiérrez, Olivier Parcollet, Nikolay Prokof’ev, and Andreas Weichselbaum for inspiring discussions and the authors of [20] and [63] for sharing their data. O.G. is supported by the NSF under Grants No. PHY-2112738 and PHY-2328774. M.G. is supported by by the Israel Science Foundation (ISF) and the Directorate for Defense Research and Development (DDR&D) grant No. 3427/21, by the ISF grant No. 1113/23 and by the US-Israel Binational Science Foundation (BSF) Grant No. 2020072. G.C. is supported by the ISF (Grant No. 2902/21) and by the PAZY foundation (Grant No. 318/78). We acknowledge high-performance computing support of the R2 compute cluster (DOI: 10.18122/B2S41H) provided by Boise State University’s Research Computing Department, the chimera cluster provided by UMass Boston Research Computing, and the Unity cluster.
@@ Appendix A Numerical benchmarks
We have performed an extensive array of tests to ensure the validity of our numerical results across the range of physical parameters considered. For internal consistency, we employed two independent diagrammatic expansions: the spin–bath coupling expansion and the diabatic coupling expansion combined with a cumulant inchworm expansion. We also varied all auxiliary numerical parameters, such as time step and maximal expansion order, to exclude potential truncation errors. In the Ohmic regime, we reproduce the analytically known Toulouse point and ensure consistency with numerical data from previous work with the inchworm algorithm [41, 42], which was in turn benchmarked with QUAPI and HEOM.
Figure 5 compares our data at very low temperature ($\beta\Delta=100$) to the numerically exact results obtained at zero temperature using ML-MCTDH [63]. Our dynamics data also agree with the data for the decoupled initial condition shown in Fig. 5 of Ref. [46].
Figure 5: Time-evolution of $\sigma_{z}(t)$ for $\omega_{c}=5\Delta$ and $\alpha=0.2$ for different values of $s$. The black lines correspond to data from Fig. 10(a) of [63] (note that units were adjusted to match our convention for the model Hamiltonian). They agree well with the inchworm QMC data from this work (color-coded) for both the spin-bath coupling expansion (left panel) and diabatic cumulant expansion (right panel).

### v2

@@ Abstract
We investigate the real-time dynamics of the sub-Ohmic spin–boson model across a broad range of coupling strengths, using the numerically exact inchworm quantum Monte Carlo algorithm. From short- and intermediate-time dynamics starting from an initially decoupled state, we extract signatures of the zero-temperature quantum phase transition between localized and delocalized states. We show that the dynamical phase diagram thus obtained differs from the equilibrium phase diagram in both the values of critical couplings and the associated values of the critical exponent. We also identify and quantitatively analyze two competing mechanisms for the crossover between coherent oscillations and incoherent decay. Deep in the sub-Ohmic regime, the crossover is driven by the damping of the oscillation amplitude, while closer to the Ohmic regime the oscillation frequency itself drops sharply to zero at large coupling.
@@ Introduction.
Although at finite $\omega_{c}$ delocalization is possible even at zero temperature, it can take a very long time for the system to delocalize [1]. Studies of the QPT have thus so far been performed directly in equilibrium [16, 17, 18, 19, 20, 21, 22, 23] or in the infinite time limit [35], even though experimental probing of the dynamics can only access short- and intermediate-time properties. Another important feature manifested by the transient dynamics is the coherence of the decay process [46, 47, 35, 48, 49], which is always probed in terms of short-time dynamics. Usually, the change from well-defined oscillations at weak spin–bath coupling to incoherent decay at strong coupling is characterized as a crossover [46, 47, 48, 49].
In this Letter, we extract signatures of the localization transition directly from the dynamics, showing that the result differs dramatically from the localization transition in equilibrium for $s\gtrsim 0.4$. We also reveal two distinct mechanisms that drive the change from coherent to incoherent decay: Overdamping of the oscillation amplitude that occurs at all values $10<s<$ 1; and a sharp decrease in oscillation frequency that is only observed for $s\gtrsim 0.5$. While the former is indeed a smooth crossover, the latter mechanism has the hallmarks of a sudden transition.
@@
The transient behavior of the sub-Ohmic spin–boson model depends strongly on the bath initial condition [46]. One usually considers the “decoupled” initial condition (where the bath is decoupled from the spin subsystem) or the “shifted” initial condition (where the bath is at equilibrium with the spin subsystem state fixed) [49]. Here, the initial system is assumed to be decoupled and the total density matrix can be represented by the factorized state $\rho_{0}=\rho_{\mathrm{s}}\otimes\rho_{\mathrm{b}}$, where the spin subsystem is in state $1$ ($\rho_{\mathrm{s}}=|1\rangle\langle 1|$) and the bath in thermal equilibrium, $\rho_{\mathrm{b}}=e^{-\beta H_{\mathrm{b}}}/\textnormal{Tr}\left\{e^{-\beta H_{\mathrm{b}}}\right\}$. Following the initial preparation, we focus on the dynamics of the population difference between the two spin states, $\langle\hat{\sigma}_{z}(t)\rangle=\textnormal{Tr}\left\{\rho_{0}e^{iHt}\hat{\sigma}_{z}e^{-iHt}\right\}$.
@@
To obtain numerically exact dynamics of $\langle\hat{\sigma}_{z}(t)\rangle$, we employ a real-time variant [50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59] of the continuous-time QMC algorithm [60, 61, 62]. To bypass the dynamical sign problem, which makes it exponentially difficult to access long times, we implemented an “inchworm” algorithm [40]. The idea behind this algorithm is that short-time propagators are less expensive to calculate than long-time propagators, and can be recycled to construct easier expansions for propagators over longer timescales in subsequent Monte Carlo steps. The inchworm algorithm has been successfully applied to study the dynamics of the spin–boson model with the Debye spectral density [41, 42], in which two types of diagrammatic expansions were developed: the spin–bath coupling expansion and the diabatic coupling expansion (the latter is combined with a cumulant inchworm expansion). The results are consistent, although in different parameter regimes one of the expansions may be more efficient than the other [42]. In the context of the Ohmic and sub-Ohmic spin–boson model, we validated the accuracy of our inchworm Monte Carlo results by detailed comparisons with numerically exact results that are available at zero temperature using ML-MCTDH [63], see Supplemental Material (SM) [64] for more details.
@@ Results.
We only consider unbiased systems, i.e., $\epsilon=0$. We set $\omega_{c}=10\Delta$ throughout and use $\Delta$ as the unit of frequency. We have found the following inverse temperatures to be indistinguishable from the zero temperature limit within the statistical errors of the simulation: $\beta\Delta=100$ for $0.50.2\leq s<$ 0.5, and $\beta\Delta=20$ for $0.5\leq s\leq 1$.
In Fig. 1 we qualitatively compare the time-evolution of $\langle\sigma_{z}(t)\rangle$ as a function of coupling $\alpha$ in the deep sub-Ohmic ($s=0.2$) and near-Ohmic ($s=0.8$) regimes. In both regimes, the system undergoes a transition from a delocalized state at weak coupling to a localized state at strong coupling. The corresponding critical coupling $\alpha^{*}$, as obtained by criterion (a), increases with increasing $s$. We also observe the crossover between coherent oscillations at weak coupling and incoherent decay at strong coupling. In the deep sub-Ohmic regime, the crossover is damping-driven as the oscillation frequency remains essentially unchanged and the amplitude is more strongly damped when $\alpha$ increases (criterion (c)). In the near-Ohmic regime on the other hand, in addition to the decrease in amplitude, the oscillation frequency also decreases with increasing $\alpha$ and completely vanishes at large $\alpha$ (criterion (b)). The $\alpha$ dependence of the renormalized oscillation frequency can be captured qualitatively by a theoretical treatment based on the analytical renormalization group method used in Ref. [65]; the result is displayed as the dashed black curves representing oscillation peaks in Fig. 1 (see SM [64] for further discussion).
A quantitative analysis of the fit parameters obtained by fitting the numerical data to Eq. (4) sheds light on the nature of the localization/delocalization transition and the coherence/incoherence crossover. To distinguish between the localized and delocalized phase we examine the offset coefficient $c$ of the fit, shown in Fig. 2. A sharp transition from zero to finite $c$ is observed for $s\lesssim 0.5$. For larger $s$ (shown in Fig. S5 of the SM [64]), the offset $c$ increases more smoothly with $\alpha$, resulting in a larger uncertainty on the critical value $\alpha^{*}$. For the equilibrium QPT the critical exponents of the transition are known [2, 13, 14, 15]. At $s\leq 0.5$ the exponent $\beta$ defined via $\langle\sigma_{z}\rangle\propto(\alpha-\alpha_{c})^{\beta}$ equals $1/2$, implying that the offset is a concave function of $\alpha$ [22]. In contrast, the corresponding scaling of the transient phase diagram analyzed in this work indicates a scaling exponent above 1, as the offset is a convex function of $\alpha$, see Fig. 2 and SM [64] for more information.
@@
Next, to characterize the coherence/incoherence crossover, we focus on the $\alpha$-dependence of the renormalized frequency $\Omega/\Delta$, criterion (b), and its ratio to the damping coefficient $\Omega/\gamma_{1}$, criterion (c), both shown in Fig. 3. Interestingly, after a small initial dip, the frequency increases with increasing $\alpha$ for small values of $s$ ($s<0.5$). In contrast, for $s\gtrsim 0.5$ the frequency monotonically decreases with $\alpha$ and eventually sharply drops to zero at sufficiently strong coupling. This sharp frequency drop is not predicted by the renormalization group treatment [65, 64] and is distinct from the localization/delocalization transition. The widely accepted notion is that the crossover between coherent and incoherent dynamics is smooth. As opposed to the sharp frequency-driven incoherence transition, criterion (c) identifies the standard hallmark of a smooth incoherence crossover, which is driven by the over-damping of the oscillation. The $\Omega/\gamma_{1}$ ratio decreases with increasing $\alpha$ for all values of $s$ and we choose the (arbitrary) threshold $\Omega/\gamma_{1}=1$ to indicate the over-damping crossover. The two incoherence mechanisms are clearly distinct from each other, both in terms of their sharpness and in terms of the parameter range in which they were observed to occur.
@@
Figure 3: Two distinct hallmarks of the coherence/incoherence crossover: the renormalized oscillation frequency, $\Omega/\Delta$, (top panel) and the ratio between $\Omega$ and the corresponding damping coefficient $\gamma_{1}$ (bottom panel). The inset in the top panel magnifies the region where the frequency behaves non-monotonically for small values of $s$. The horizontal line in the bottom panel marks the $\Omega/\gamma_{1}=1$ threshold.
@@
A summary of our quantitative results is presented through the phase diagram in the parameter space of $\alpha$ and $s$, see Fig. 4. The transient critical coupling $\alpha^{*}$ distinguishing the localized and delocalized phases agrees well with the $\alpha_{c}$ of the equilibrium QPT for a model with sudden frequency cutoff [66, 20, 67] for sufficiently small $s$, where the critical couplings are also well-described by the analytical prediction from the generalized polaron ansatz [68]. For $s\gtrsim 0.4$ the two values begin to deviate. The frequency- and the damping-driven crossover lines depicted in Fig. 4 are clearly distinct from each other, with the frequency-driven crossover only observed to occur in the near-Ohmic regime, while the damping-driven crossover can occur at any value of $s$ (though it is numerically challenging to observe at small $s$). Coherent and incoherent regimes are observed on either side of the localization/delocalization line. For the Ohmic spin–boson model the location of the incoherence transition is known analytically to occur at $\alpha=0.5$ (Toulouse point) [2]. Our numerical results for the frequency-driven transition (criterion (b)) are consistent with the Toulouse point. The $\Omega=\gamma_{1}$ line of the damping-driven criterion (c) yields a different value for the Ohmic transition point.
@@
Figure 4: Transient dynamical phase diagram of the sub-Ohmic spin–boson model. The red shaded region is our numerical result for $\alpha^{*}$ that characterizes the dynamical localization-delocalization transition (criterion (a)), with the area marking a confidence interval. For $s\lesssim 0.4$ it agrees well with the equilibrium $\alpha_{c}$ from a QMC calculation for a model with sudden cutoff [20] (red circles and solid line) [69]. The dotted and dashed lines are the two characterizations of the change from coherent to incoherent dynamics, corresponding to criteria (b) and (c), respectively. The system tends toward incoherence above the lines, i.e. at higher values of $\alpha$. The frequency-driven crossover (blue dotted line) was only observed in the near-Ohmic regime, for the range of parameters considered. The damping-driven crossover (green dashed line), which we defined as $\gamma_{1}=\Omega$, was observed in a wider parameter range.
@@ Conclusions.
The inchworm algorithm used here is efficient over a wide range of parameters. Looking forward, it can also be used to explore dynamics and full charge/energy counting statistics at higher temperatures [70, 71, 72, 73], in more general models [74], or in the presence of intrinsically nonequilibrium drives such as multiple baths at different thermodynamic parameters. Extensions exist for investigating dynamics over very long timescales [75] and nonequilibrium steady states [76, 77]. On a more general note, our work points the way toward the ability to investigate and potentially control transient “phase diagrams” in numerous scenarios, opening up intriguing prospects for revealing new physics in experiments, theory and simulations.
Acknowledgments.
We thank Jianshu Cao, Jan von Delft, Carlos González-Gutiérrez, Olivier Parcollet, Nikolay Prokof’ev, and Andreas Weichselbaum for inspiring discussions and the authors of Refs. [20] and [63] for sharing their data. O.G. is supported by the NSF under Grants No. PHY-2112738 and No. PHY-2328774. M.G. is supported by the Israel Science Foundation (ISF) and the Directorate for Defense Research and Development (DDR&D) Grant No. 3427/21, by the ISF grant No. 1113/23 and by the U.S.-Israel Binational Science Foundation (BSF) Grant No. 2020072. G.C. is supported by the ISF (Grant No. 2902/21) and by the PAZY foundation (Grant No. 318/78). We acknowledge high-performance computing support of the R2 compute cluster (DOI: 10.18122/B2S41H) provided by Boise State University’s Research Computing Department, the chimera cluster provided by UMass Boston Research Computing, and the Unity cluster.
@@ Appendix A Numerical benchmarks
We have performed an extensive array of tests to ensure the validity of our numerical results across the range of physical parameters considered.
For internal consistency, we employed two independent diagrammatic expansions: the spin–bath coupling expansion and the diabatic coupling expansion combined with a cumulant inchworm expansion. We also varied all auxiliary numerical parameters, such as time step and maximal expansion order, to exclude potential truncation errors. For the spin-bath coupling expansion a maximal order of 4 was found to be sufficient to achieve convergence and to produce accurate results. For the cumulant expansion the maximal order was set to 14, but in practice, such high orders were never generated by the simulations, so that effectively the maximal order can be considered to be infinite.
The maximal times simulated varied from $t_{max}\Delta=20$ to $t_{max}\Delta=80$, depending on the dynamical behavior of the system. Larger values of $t_{max}$ were needed in strongly incoherent regimes that delocalize slowly (large $s$ and large $\alpha$), as well as in strongly coherent regimes with small damping (small $s$ and small $\alpha$). We checked consistency of our fit results for different values of $t_{max}$ for representative physical parameter sets.
In the Ohmic regime, we reproduce the analytically known Toulouse point and ensure consistency with numerical data from previous work with the inchworm algorithm [41, 42], which was in turn benchmarked with QUAPI and HEOM.
Figure S1 compares our data at very low temperature ($\beta\Delta=100$) to the numerically exact results obtained at zero temperature using ML-MCTDH [63]. Our dynamics data also agree with the data for the decoupled initial condition shown in Fig. 5 of Ref. [46].
Figure S1: Time-evolution of $\sigma_{z}(t)$ for $\omega_{c}=5\Delta$ and $\alpha=0.2$ for different values of $s$. The black lines correspond to data from Fig. 10(a) of [63] (note that units were adjusted to match our convention for the model Hamiltonian). They agree well with the inchworm QMC data from this work (color-coded) for both the spin-bath coupling expansion (left panel) and diabatic cumulant expansion (right panel).